Today Rana bolted down the cavity mirrors and did some alignment. I measured the distances between mirrors and have worked out a rough mode matching scheme. I'll post the beamscan measurements separately since they will also be useful for the final mode matching calculation for the astigmatism in the cavity (and beam). We probably want to make a beam measurement for the second path as well after it comes out of the AOM double-pass loop.
I've put the lenses in place for the rough mode matching and started aligning the cavity again. The astigmatism in the beam is clearly worse than it was before because of the 6m radius curved mirrors.
Edit --- I've spent a while trying to get the mode aligned. It still looks pretty messy though there is a 00 trying to break through every now and again. I'm not sure why it looks so screwy - the light around the edges just seems to be in random patterns which we didn't see before. I wonder if this is just light being lost because of the astigmatism?